# Eye-tracking-Driven Shared Control for Robotic Arms: Wizard of Oz Studies to Assess Design Choices

### v1

@@ 6. Limitations
Additional questions: The safety concerns were not reflected in the responses to the online survey. Including questions such as ”Would you feel safe interacting with the robot?” would provide a more comprehensive understanding of this topic. However, other studies show that potential users are more likely to accept the risks of using the system than to reject it and become dependent on caregivers or family members (Baumeister et al., 2021). This can also be seen here, as the robot-experienced group is aware of the risks of human-robot interaction but is not fearful of interacting with the robot. Specific Wizard of Oz design: The Wizard of Oz system is an eye-tracking-driven design in which objects are selected without the need for a GUI. This approach is uncommon since prototype systems face challenges such as interpreting the desired task, error estimation due to head movements, and a lack of feedback modalities (Fischer-Janzen et al., 2024). Researchers using these adaptation suggestions should check whether the findings can be applied to different input designs. Generalized approach: This approach was not tailored to specific disabilities, so it may lack details necessary to adapt to individual users’ needs. Future work can use this generalization to more easily locate these demands.
@@ 7. Conclusion
The goal of this work was to obtain feedback on an eye-tracking-driven shared control for Assistive Robotic Arms to identify community expectations, desired tasks, and technical challenges. The factors for a successful implementation of a shared control were identified and presented. A Wizard of Oz design was created prior to prototyping, as suggested by other researchers.
Based on this design, two studies were conducted. An online survey designed for people with disabilities, their families and friends, and caregivers provided feedback on the community’s expectations for the system, as well as the desired task the robot should be able to perform. A hands-on study with participants with and without robotics experience provided insight into the technical requirements and feedback differences between the video presentation available for the online survey and the real-world application. Gaze data was recorded to measure dwell times and influence dependencies related to the eye-tracking control.
Dependencies for the design of eye-tracking-driven robot controls were uncovered, ranging from user demands on task prioritization, to accuracy limiting factors in eye-tracking. As a consequence, improvements to the initial idea were presented to facilitate the usage of the system.
@@ Acknowledgements.
Both studies were approved by the ethics commission of the Offenburg university of Applied Sciences. The study designs were checked to comply to the ACM Publications Policy on Research Involving Human Participants and Subjects. The Tobii Pro Glasses 3 was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number INST898/33-1. Further, we would like to thank Atikkhan Faridkhan Nilgar and Karl Tschurtschenthaler for the great discussions in statistics and eye-tracking evaluation related to this topic. We would like to thank Katrin-Misel Ponomarjova to proof read the translations of the questionnaire. Lastly, we want to thank Kim Zähringer for her assistance with the robot and video making.
Appendix A Participant statements of hands-on study
@@
24 Exp. Fork was recognized well, even though you looked briefly at the bottle during the time you had to aim. The pause between aiming and starting the pick and place task was too long!

### v2

@@ 6. Limitations
Online study: The qualitative analysis of the online survey provided limited insight into the participants’ everyday lives, meaning the data cannot be generalized to a larger population. To counter this issue, we enriched and compared our information with suitable literature. For future studies, however, we have decided to conduct interviews with participants to minimize potential hurdles due to a lack of assistive devices or display problems on end devices. Q4 was tailored to evaluate whether hurdles exist. No indications were found. Additionally, no emails stating technical hurdles were received from medical professionals, family, or PSMD. The online survey might have benefited from additional questions. The safety concerns were not reflected in the survey responses. Including questions such as ”Would you feel safe interacting with the robot?” would have provided a more comprehensive understanding of this topic. However, other studies show that potential users are more likely to accept the risks of using the system than reject it, becoming dependent on medical professionals or family members (Baumeister et al., 2021). This can also be seen here, as the group with robot experience is aware of the risks of human-robot interaction, but is not fearful of interacting with the robot. The survey did not ask about the occupation of the medical professionals. Therefore, opinions may differ depending on the occupational field (e.g., doctors, caregivers, and nurses). The online survey was designed for quantitative analysis. Due to the limited number of participants, a qualitative analysis was performed. Therefore, not all answers were discussed in detail, since they are not representative (e.g., Q5 rel./fr.: time spent on care). These results can be accessed in the supplementary materials. As discussed in Section 3.3, participants did not interact directly with the robot, which might affect perception.
Specific Wizard of Oz design: The Wizard of Oz system does not vary in intent interpretation due to its hard-coded trajectories. Variable behavior was excluded from this study, since comparing outcomes between participants would not be applicable. However, varying behavior in shared control changes the perception of the system, as shown in (Aronson et al., 2018). We presented a design driven by eye-tracking, in which objects are selected without the need for a GUI. This approach is uncommon, since prototype systems face challenges such as interpreting the desired task, estimating errors due to head movements, and lacking feedback modalities (Fischer-Janzen et al., 2024). If differences in control design exist in relation to the WoZ design, the results may not be transferable to other controls.
Generalized approach: Since this approach was not tailored to specific disabilities, it may lack the details necessary to adapt to the needs of individual users. Future work can use this generalization to more easily identify these requirements.
Transfer to other eye trackers: Only the Tobii Pro Glasses 3 eye tracker was used, which is more accurate than other eye-tracking glasses on the market. However, the effect on the bounding boxes in object detection can vary greatly depending on the model used. Due to the scope of this paper, the test was not reproduced with other eye trackers.
@@ 7. Conclusion
The goal of this work was to obtain feedback on an eye-tracking-driven shared control for assistive robotic arms in order to identify community expectations, desired tasks, and technical challenges. While previous studies offer valuable findings, they vary between robot types and involve alternative stakeholder groups. As a result, the available insights are generally only limited transferable and not specific to the control for an assistive robotic arm examined here. As suggested by other researchers, a Wizard of Oz design was created prior to prototyping. The factors necessary for successfully implementing shared control were identified and presented.
Based on this WoZ design, two studies were conducted. An online survey designed for PSMD, their families and friends, and medical professionals. This survey provided feedback on the community’s expectations for the system and the tasks the robot should be able to perform. A hands-on study with participants with and without robotics experience provided insight into the technical requirements and feedback differences between the video presentation available for the online survey and the real-world application. Gaze data was recorded to measure dwell times and influence dependencies related to the eye-tracking controls.
The duality of both studies provided insight into the usability and functionality of the system. Task prioritization was contextualized with current literature and approaches that researchers presented to realize shared controls. Dependencies were uncovered for the design of eye-tracking-driven robot controls, such as the influence of object size, geometry, gaze location, and user intent on accuracy. The impact on object detection or online task intent recognition based approaches were assessed and discussed. Consequently, improvements and open challenges related to the initial idea were presented.
@@ Acknowledgements.
Both studies were approved by the ethics commission of the Offenburg University of Applied Sciences. The study designs were checked to comply to the ACM Publications Policy on Research Involving Human Participants and Subjects. The Tobii Pro Glasses 3 were funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project number INST898/33-1. Further, we would like to thank Atikkhan Faridkhan Nilgar and Karl Tschurtschenthaler for the great discussions in statistics and eye-tracking evaluation related to this topic. We would like to thank Katrin-Misel Ponomarjova for proofreading the translations of the survey. We want to thank Kim Zähringer for her assistance with the robot and video making.
Appendix A Overview of Questions asked in the Online Survey
In the following an Overview of the questions asked to the participants in the online Survey are presented. They represent three versions of the survey which would be handed to the individual participant group PSMD, family and friends and medical professionals. All related answers can be found in the supplementary materials.
Table 5. This table presents all questions asked in the online survey. Questions spanning over all three columns indicate questions that were asked to each participant.
Indicator Questions asked to group
PSMD Relatives and friends Medical professionals
Q1 What age group are you in?
Q2 What gender are you?
Q3 To which of the groups below do you feel you belong?
Q4 Do you have assistance from another person in completing this questionnaire? Did the person you care for participate in this survey? Please estimate the percentage of your patients or clients who have movement difficulties in their arms?
Q5 If so, what is the reason you need support? How much time do you spend on care each day (in hours)? Let’s assume that an assistance system could take over all daily tasks. Which tasks would you most like to hand over in order to facilitate care? Name 1-3 favorites.
Q6 How often does your impairment affect your daily life? Do you already use assistive devices to help you provide care? -
Q7 In terms of your physical disability, has your condition improved, worsened, or stayed the same in the last 6 months? If yes, which assistive devices do you use? -
Q8 Select the body movements that you can control without restrictions - -
Q9 Do you already use a robotic gripper arm, for example on an electric wheelchair? Does the person(s) you care for already use a robotic gripping arm, e.g. on an electric wheelchair?
Q10 If so, how do you control it? If so, how does the person control this robotic arm?
Q11 Are you satisfied with your current system? Are you satisfied with the system in terms of startup and function?
Q12 If no, indicate why you are not satisfied with the system. If no, why are you not satisfied with the system?
Q13 In relation to your disability, would it be possible for you to use the system presented in the video every day? In which tasks would you or the person you care for most like to be supported by this system in daily care? Please name your 1-3 favorites.
Q14 If no, what changes would you suggest to make the system easier for you to use? Would it be possible for the person you care for to use the system presented in the video every day?
Q15 In which tasks would you most like to be supported by this system? Please name your 1-3 favorites. If no, what changes would you suggest to make the system easier to use for you and the person you care for?
Q16 Do you have any other comments on the questions asked or the system?
Q17 Would you like to have had more detailed information on how to use the system to answer the questions asked here?
Q18 If yes, what information would you have liked to have?
Appendix B Participant statements of hands-on study
@@
24 Exp. Fork was recognized well, even though you looked briefly at the bottle during the time you had to aim. The pause between aiming and starting the pick and place task was too long.